# Flat Space Holography via AdS/BCFT

### v1

@@ 1 Introduction
The purpose of this paper is to provide a new approach to the flat space holography. In particular, we would like to consider a method to embed it in the framework of AdS/CFT. This is because it allows us to understand the flat space holography as a version of AdS/CFT, which is quite established and well understood. For this purpose, we employ the AdS/BCFT correspondence Takayanagi:2011zk ; Fujita:2011fp ; Karch:2000gx , which argues that the gravity dual of a CFT on a $d$ dimensional manifold $N_{d}$ with a conformal boundary (BCFT) is dual to a part of AdS geometry $M_{d+1}$ which is surrounded by the manifold $N_{d}$ and a surface $Q_{d}$ called the end-of-the-world brane (EOW brane) such that $\partial M_{d+1}=N_{d}\cup Q_{d}$. One advantage of the AdS/BCFT is that we can extend the holographic calculation of entanglement entropy Ryu:2006bv ; Ryu:2006ef ; Hubeny:2007xt to the setups with EOW branes by simply allowing the extremal surface can end on EOW branes Takayanagi:2011zk ; Fujita:2011fp . Another advantage is that it can have the third dual interpretation by regarding the CFT dual as a $d$ dimensional CFT coupled to a $d$ dimensional gravity on the EOW brane via the brane world holography Karch:2000ct ; Gubser:1999vj ; Randall:1999ee ; Randall:1999vf .
@@
The EOW branes, the most important ingredient of the AdS/BCFT, are characterized by values of tension $T$. Depending on values of tension, they are classified into three classes such that the induced metric on a given EOW brane becomes either AdS, dS or flat space. The most standard one is the AdS brane, which has successfully been applied to the black hole information problem Almheiri:2019hni ; Almheiri:2019psf as the holographic entanglement in the AdS/BCFT can explain the island formula Penington:2019npb ; Almheiri:2019psf , which was explicitly confirmed in Suzuki:2022xwv ; Izumi:2022opi . For applications to cosmological models refer to Cooper:2018cmb ; Antonini:2019qkt ; VanRaamsdonk:2021qgv ; Omiya:2021olc ; Antonini:2022blk ; Waddell:2022fbn . The dS branes have not been completely well-understood because they are exotic from the viewpoint of CFTs as they are dual to space-like boundaries in CFTs Akal:2020wfl ; Chen:2020tes ; Hao:2024nhd . Nevertheless, several interesting applications have been considered recently, including the gravity dual of cross cap states Wei:2024zez and an embedding dS/CFT into AdS/BCFT Fujiki:2025yyf .
@@
The least studied EOW branes are the flat branes which are realized for special values of tension. Via the brane-world holography it is directly related to holography for flat spacetimes. Thus in the present paper we would like to explore properties of flat EOW branes in order to get insights on how the holography for flat spacetimes looks like. The AdS/BCFT approach allows us to compute various quantities such as entanglement entropy and correlation functions as we will show in this paper. Interestingly, flat EOW branes are dual to null boundaries in CFTs. Such boundaries in field theories have not been well-understood and are worth exploring by combining the field theoretic and holographic calculations.
@@
Using a flat EOW brane, we will consider three setups, called type I, II and III (refer to Fig.3). In the type I and II model, the brane has a negative and positive tension, respectively. In the type I case, the CFT is defined on a diamond with null boundaries, where the boundaries play a role similar to the final state projections. On the other hand, in the type II case, the CFT on the diamond is coupled to flat space gravity through the null boundaries. In the type II construction, we will find that our AdS/BCFT consideration reproduces the Flat/Carrollian CFT (CCFT) correspondence, where the Carrolian CFT is realized on the null boundary of the flat EOW brane. We can pick up purely the flat space gravity by sandwiching the bulk AdS region by two flat EOW branes and by applying the wedge holography Akal:2020wfl , which is called the type III setup and can make the above argument clearer.
@@ 2.1 Flat EOW brane
The main purpose of this paper is to examine the second case (ii), which has not been well-studied in the context of holography till now. If we employ the brane-world holography idea to the AdS/BCFT, we can obtain the third interpretation in terms of the CFT on the $d$ dimensional manifold coupled to a $d$ dimensional gravity on the EOW brane along the boundaries, which is called the double holography and has been often employed to analyze the black hole information problems. If we apply this interpretation, we can regard the setup (ii) as a CFT coupled to a gravity on a flat spacetime. Therefore, we can embed the problem of the holography in flat spacetime, which is far from well-understood even now, into the AdS/BCFT setups, which is much more controllable.
@@ 2.4 Type II Setup
Instead, in the type II setup, we pick up the interior region of the flat EOW brane with $T=\frac{d-1}{R}$, as depicted in the middle left panel of Fig.3. This region can be infinitely extended in both the future and past direction $\tau\to\pm\infty$, which creates the AdS boundary region in these directions. Therefore, its CFT dual looks like the CFT which surrounds the diamond (2.9) and its null boundaries. Via the double holography, this is equivalent to the CFT coupled to the flat space gravity on the diamond as depicted in the middle center panel of Fig.3. Hence, this provides an interesting model which includes the flat space holography.
@@ 2.5 Type III Setup
In the type II setup, the field theory dual includes the large ‘bath’ theory, which extends outside of the diamond. To extract the flat space holography, we can remove the large bath by adding one more flat EOW brane and by focusing on the region between the two EOW branes, sketched in the bottom panels of Fig.3. This consideration leads to the type III setup. The AdS boundary of this model only consists of the null edges of the diamond. Thus, its CFT dual is given by a d1d-1-dimensional Carollian CFT (CCFTd-1) which lives on the null lines $|\tau+\phi|=|\tau-\phi|=\pi$. Note that the insertion of the flat EOW brane breaks the isometry of the AdSd+1 with the remaining symmetry group ISO(1,d1)ISO(1,d-1), which is also the global symmetry group of the CCFTd-1. Via the double holography, we can also regard this as the flat space gravity on the diamond. Thus, this provides an interesting setup of flat space holography.
@@ 3.2 Holographic entanglement entropy in type II setup
$D_{13}=\log\big(\cos(\tau_{1}-\tau_{3})\cosh\rho_{3}-\cos(\phi_{1}-\phi_{3})\sinh\rho_{3}\big{missing})-\log\epsilon,$ (3.20)
@@ 3.3 Holographic entanglement entropy in type III setup
Compared to the entanglement entropy in CCFT2 Bagchi:2014iea ; Jiang:2017ecm ; Apolo:2020bld
@@ 4 Holographic correlation functions
Next we would like to turn to holographic correlation functions to further explore physical properties. Basically, we can generalize the computations of correlation functions based on the bulk-boundary relation Gubser:1998bc ; Witten:1998qj to our setups with flat EOW branes. We will focus on the two-point functions and one-point functions.

### v2

@@ 1 Introduction
The purpose of this paper is to provide a new approach to the flat space holography. In particular, we would like to consider a method to embed it in the framework of AdS/CFT. This is because it allows us to understand the flat space holography as a version of AdS/CFT, which is quite established and well understood. For this purpose, we employ the AdS/BCFT correspondence Takayanagi:2011zk ; Fujita:2011fp ; Karch:2000gx , which argues that the gravity dual of a CFT on a $d$ dimensional manifold $N_{d}$ with a conformal boundary (BCFT) is dual to a part of AdS geometry $M_{d+1}$ which is surrounded by the manifold $N_{d}$ and a surface $Q_{d}$ called the end-of-the-world brane (EOW brane) such that $\partial M_{d+1}=N_{d}\cup Q_{d}$.
One advantage of the AdS/BCFT is that we can extend the holographic calculation of entanglement entropy Ryu:2006bv ; Ryu:2006ef ; Hubeny:2007xt to the setups with EOW branes by simply allowing the extremal surface can end on EOW branes Takayanagi:2011zk ; Fujita:2011fp . Another advantage is that we can identify the setups of EOW branes in AdS geometries with the brane world models, namely Randall-Sundrum models, so called RS1 and RS2 model, Randall:1999ee ; Randall:1999vf . It has been known for a while that the brane world models enjoy an intriguing holographic duality, which generalizes the AdS/CFT Gubser:1999vj ; Giddings:2000mu ; Arkani-Hamed:2000ijo ; Rattazzi:2000hs ; Perez-Victoria:2001lex ; Soda:2010si . This brane world holography argues that the classical gravity which lives the $d+1$ dimensional AdS bulk surrounded by the brane(s) is dual to a system where a $d$ dimensional CFT with a finite UV cut off is coupled to the $d$ dimensional dynamical quantum gravity. The gravity on the brane is dynamical because we impose the Neumann boundary on it. Notice that the $d$ dimensional gravity on the brane is known to be described by a modification of Einstein gravity Garriga:1999yh ; Shiromizu:1999wj ; Gubser:1999vj ; Giddings:2000mu and this modification is dual to the loop corrections to the graviton propagator due to the exchange of the CFT matter via the holography. This brane world holography adds the third dual interpretation to the AdS/BCFT duality, which we will fully employ in this paper.
@@
The EOW branes, the most important ingredient of the AdS/BCFT, are characterized by values of tension $T$. Depending on values of $T$, they are classified into three classes such that the induced metric on a given EOW brane becomes either AdS, dS or flat space. The standard one is the AdS brane, where gravity localizes on the brane Karch:2000ct , called Karch-Randall model, and thus describe quantum gravity on AdS. This has successfully been applied to the black hole information problem Almheiri:2019hni ; Almheiri:2019psf as the holographic entanglement in the AdS/BCFT can explain the island formula Penington:2019npb ; Almheiri:2019psf ; Chen:2020uac , which was explicitly confirmed in Suzuki:2022xwv ; Izumi:2022opi . For applications to cosmological models which employ the AdS/BCFT holography, refer to Cooper:2018cmb ; Antonini:2019qkt ; VanRaamsdonk:2021qgv ; Omiya:2021olc ; Antonini:2022blk ; Waddell:2022fbn . The holographic interpretations of dS branes have not been completely well-understood because they are exotic from the viewpoint of CFTs as they are dual to space-like boundaries in CFTs Akal:2020wfl ; Chen:2020tes ; Hao:2024nhd . Nevertheless, several interesting applications have been considered recently, including the gravity dual of cross cap states Wei:2024zez and an embedding dS/CFT into AdS/BCFT Fujiki:2025yyf .
@@
From the AdS/BCFT viewpoint, the least studied EOW branes are the flat branes which are realized for special values of tension. However, we know that the flat space EOW brane leads to a holography for flat spacetimes, where the $d$ dimensional quantum gravity coupled with a CFT on a flat space gets equivalent to the $d+1$ dimensional bulk AdS surrounded by the brane(s), owing to the much earlier and many works on the brane-world holography, including Gubser:1999vj ; Giddings:2000mu ; Arkani-Hamed:2000ijo ; Rattazzi:2000hs ; Perez-Victoria:2001lex ; Soda:2010si . Motivated by this, the main purpose of this paper is explore the implications of holography for the flat EOW branes by combining the ideas of the modern AdS/BCFT with the celebrated brane-world holography in order to get insights on how the holography for flat spacetimes looks like. The main difference between our AdS/BCFT approach and the original brane-world holography is that in the former, the dual CFT always live on the UV boundary and the EOW brane is treated as the boundary condition in the IR, while in the latter, the finite cut off CFT lives on the brane. Our AdS/BCFT approach has the advantage that we can compute various quantities such as entanglement entropy and correlation functions in the fine grained CFT situated at the UV boundary of AdS, as we will see in this paper. Another intriguing point is that flat EOW branes are dual to null boundaries in CFTs. Such boundaries in field theories have not been well-understood till now and are worth exploring by combining the field theoretic and holographic calculations.
@@
Using a flat EOW brane, we will consider three setups of AdS/BCFT, called type I, II and III (refer to Fig.3). In the type I and II model, the brane has a negative and positive tension, respectively. In the type I case, the CFT is defined on a diamond with null boundaries, where the boundaries play a role similar to the final state projections. On the other hand, in the type II case, the CFT on the diamond is coupled to flat space gravity through the null boundaries. In the type II construction, we will find that our AdS/BCFT consideration reproduces the Flat/Carrollian CFT (CCFT) correspondence, where the Carrolian CFT is realized on the null boundary of the flat EOW brane. We can pick up purely the flat space gravity by sandwiching the bulk AdS region by two flat EOW branes and by applying the wedge holography Akal:2020wfl , which is called the type III setup and can make the above argument clearer. We would like to note that in the Poincare AdS coordinate patch, these three setups are not new and have been discussed by many authors. The type I model corresponds to the hard wall model with an IR cut off, which has often been used in AdS/QCD Erlich:2005qh . The type II model and type III model are equivalent to the RS2 model Randall:1999vf and RS1 model Randall:1999ee , respectively. However, we would like to emphasize that in this paper we are mainly interested in the holography in the global AdS coordinate patch where we can always find the asymptotically AdS UV boundary, where we expect the dual fine grained CFT lives.
@@ 2.1 Flat EOW brane
The main purpose of this paper is to examine the second case (ii), which has not been well-studied in the context of holography till now. If we employ the brane-world holography idea to the AdS/BCFT, we can obtain the third interpretation in terms of the CFT on the $d$ dimensional manifold coupled to a $d$ dimensional gravity on the EOW brane along the boundaries, which is called the double holography and has been often employed to analyze the black hole information problems. If we apply this interpretation, we can regard the setup (ii) as a CFT coupled to a gravity on a flat spacetime. Therefore, we can embed the problem of the holography in flat spacetime, which is far from well-understood even now, into the AdS/BCFT setups, which is much more controllable. Refer also to the recent paper Fontanella:2025tbs for another different approach, where the flat space limit is taken for the whole bulk AdS space.
@@ 2.3 Type I Setup
If we regard the EOW brane as a hard wall, the type I model looks similar to the AdS/QCD model with a hard wall e.g. Erlich:2005qh . In this context the brane can be regarded as the IR cut off.
@@ 2.4 Type II Setup
Instead, in the type II setup, we pick up the interior region of the flat EOW brane with $T=\frac{d-1}{R}$, as depicted in the middle left panel of Fig.3. This region can be infinitely extended in both the future and past direction $\tau\to\pm\infty$, which creates the AdS boundary region in these directions. Therefore, its CFT dual looks like the CFT which surrounds the diamond (2.9) and its null boundaries. Via the double holography, this is equivalent to the CFT coupled to the flat space gravity on the diamond as depicted in the middle center panel of Fig.3. Hence, this provides an interesting model which includes the flat space holography. Note also that this type II setup in the Poincare patch is equivalent to the RS2 model Randall:1999vf , though in this paper we consider its global extension in order to have the asymptocially AdS UV boundary where the dual CFT lives.
@@ 2.5 Type III Setup
In the type II setup, the field theory dual includes the large ‘bath’ theory, which extends outside of the diamond. To extract the flat space holography, we can remove the large bath by adding one more flat EOW brane and by focusing on the region between the two EOW branes, sketched in the bottom panels of Fig.3. This consideration leads to the type III setup. The AdS boundary of this model only consists of the null edges of the diamond. Thus, its CFT dual is given by a d1d-1-dimensional Carollian CFT (CCFTd-1) which lives on the null lines $|\tau+\phi|=|\tau-\phi|=\pi$. Note that the insertion of the flat EOW brane breaks the isometry of the AdSd+1 with the remaining symmetry group ISO(1,d1)ISO(1,d-1), which is also the global symmetry group of the CCFTd-1. Via the double holography, we can also regard this as the flat space gravity on the diamond. Thus, this provides an interesting setup of flat space holography. Note also that this type III setup in the Poincare patch is equivalent to the RS1 model Randall:1999ee , though in this paper we consider its global extension in order to have the asymptocially AdS boundary which is codimension two.
@@ 3.2 Holographic entanglement entropy in type II setup
$D_{13}=\log\big(\cos(\tau_{1}-\tau_{3})\cosh\rho_{3}-\cos(\phi_{1}-\phi_{3})\sinh\rho_{3}\big.)-\log\epsilon,$ (3.20)
@@ 3.3 Holographic entanglement entropy in type III setup
Compared to the entanglement entropy in CCFT2 Bagchi:2014iea ; Jiang:2017ecm ; Apolo:2020bld ; Banerjee:2024ldl
@@ 4 Holographic correlation functions
Next we would like to turn to holographic correlation functions in our new AdS/BCFT setups to further explore physical properties. Basically, we can generalize the computations of correlation functions based on the bulk-boundary relation Gubser:1998bc ; Witten:1998qj to our setups with flat EOW branes. We will focus on the two-point functions and one-point functions. We would like to note that green functions in AdS geometries with branes have been worked out by many earlier papers in the context of brane world holography. For this refer to e.g. Giddings:2000mu ; Perez-Victoria:2001lex for the RS2 model and Arkani-Hamed:2000ijo ; Rattazzi:2000hs for the RS1 model, where the green functions are computed and the holographic correlation functions on the branes, dual to those in the finite cut off CFT, were found. Even though our calculations below employ essentially the same green functions, we compute the holographic correlation function not on the brane but on the UV boundary, where we expect the dual CFT lives.